# StreamMOS: Streaming Moving Object Segmentation with Multi-View Perception and Dual-Span Memory

### v1

@@ III-C Short-term Temporal Fusion
The purpose of this part is to transfer the memory feature $\mathcal{H}_{t-1}$ from the last inference to the present, so that historical spatial states of objects can be reused to guide the network to deduce object motion in the $t$ time. In this regard, we first build short-term memory bank as a bridge to store $\mathcal{H}_{t-1}$ and connect adjacent inference. Then, since $\mathcal{F}_{t}$ and $\mathcal{H}_{t-1}$ are not in the same coordinate system, we use learnable offsets [20] to adaptively find the relationship between two features and combine them by attention weight. Specifically, the $\mathcal{H}_{t-1}$ is first fed into two linear layers to produce $K$ attention weights $A_{k}$ and sampling offsets $\Delta g_{k}$. Afterward based on the offsets $\Delta g_{k}$ and coordinates $g_{k}$ of reference points in $\mathcal{F}_{t}$, a bilinear interpolation is used to gather reference values $G_{k}$ from $\mathcal{F}_{t}$. Finally, $G_{k}$ is weighted by $A_{k}$ and get a updated feature ^${\hat{\mathcal{H}}_{t}}$. The above process can be formulated as follows:
@@
where $L,K$ are the number of attention heads and reference points, respectively. The $S(\textperiodcentered\cdot\textperiodcentered)$ and $W_{l}$ represent the bilinear sampling and learnable weight of multi-head attention. Later, the ^${\hat{\mathcal{H}}_{t}}$ is processed by normalization layer and feed-forward network (FFN) to generate a renewed $\mathcal{H}_{t}$ in the current time:
@@
Here, LN is the layer normalization. And the $\mathcal{H}_{t}$ will replace $\mathcal{H}_{t-1}$ of short-term memory and participate in the prediction of segmentation decoder.
@@ III-D Reduced-Parameter Segmentation Decoder
where $\text{MLP}_{1},\text{MLP}_{2}$ mean multi-layer perceptrons that do not share weights, and $\mathcal{E}_{t}$ is the feature from point-wise encoder. According to discrete classification labels $\mathcal{O}_{t}$, we can acquire the attributes of instance, like location and size, by clustering and apply them to optimize $\mathcal{C}_{t}$ in the subsequent voting stage.
@@ III-E Long-term Voting Mechanism
Most existing approaches [4, 5, 17] focus on improving the quality of a single inference through modifying network structure. Nevertheless, in light of the inexplicability and data dependency of neural networks, this strategy may be limited. For example, for a parking car shown in Fig. 1(c), the model may predict it as stationary at this frame and dynamic at other frames. Meanwhile, due to lacking instance-level perception ability, the network may generate inconsistent results for the different parts of an object. To solve these issues, we present a voting mechanism consisting of voxel-based voting (VBV) and instance-based voting (IBV), which could be regarded as a post-processing that corrects errors in the current predicted labels $\mathcal{C}_{t}$ using historical results $\{\mathcal{M}_{t-m}\}^{M}_{m=1}$ and movable labels $\mathcal{O}_{t}$. Noted that $\mathcal{M}_{t-m}\ (m=1,...,M)$ with coordinate of $\mathcal{P}_{t-m}$ will be transformed into coordinate system of $\mathcal{P}_{t}$ to yield $\mathcal{M}^{\prime}_{t-m}$ by pose transformations $\mathcal{T}_{t-m\rightarrow t}$ in advance.
@@ III-E1 Voxel-based voting
Figure 4: The detail of our voting mechanism. It utilizes voxel-based voting (VBV) and instance-based voting (IBV) to refine coarse predictions.
Inspired by TFNet [21], we obtain the historical predictions $\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1}$ and current forecast $\mathcal{C}_{t}$ in the same coordinate system. Afterward we divide points $\mathcal{P}_{t}$ into voxels with fixed size and fill $(\mathcal{C}_{t},\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1})$ into each voxel. Then, as shown in Fig. 4(a), the most frequently predicted label acts as motion state for all points in the same voxel and incorrect labels will be updated. We brief the above procedure of VBV as: $\Omega(\mathcal{C}_{t},\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1})\mapsto\hat{\mathcal{C}_{t}}$.
@@ III-E2 Instance-based voting
Although VBV can ensure the consistency of motion states within a local area and achieve performance improvement in Tab. IV, it is difficult to achieve instance-level state unity, as shown output from VBV $\hat{\mathcal{C}_{t}}$ in Fig. 4(b). To solve this, we propose an instance-based voting based on cluster. Specifically, given predicted probability $\mathcal{O}_{t}$ from decoder, we can pick out the foreground points ${\mathcal{P}^{\prime}_{t}}$ from ${\mathcal{P}_{t}}$ and adopt DBSCAN [22] to split ${\mathcal{P}^{\prime}_{t}}$ into $S$ clusters. Then, according to the coordinates of points in each cluster, we can compute $S$ minimum 3D bounding boxes to cover all objects. Thus, we can further crop out instance-level predictions from $\hat{\mathcal{C}_{t}}$ and memory predictions $\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1}$. Finally, similar to voxel-based voting, we adopt the class label with the highest quantity as the motion state for all points in the instance and get the final prediction as: $\Phi(\hat{\mathcal{C}_{t}},\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1},\mathcal{O}_{t})\mapsto{\mathcal{M}_{t}}$.
III-E3 Long-term memory updating
We construct a memory bank with the length of $M$ to retain historical results. When a new refined prediction $\mathcal{M}_{t}$ is output from voting mechanism, we append it to memory bank and pull out the oldest result.
As a result, compared to only relying on network adaptive learning, our vote mechanism could explicitly suppress incorrect prediction and improve the consistency of segmentation by analyzing long-term prediction in voxel and instance level.
@@ III-F Loss Functions
To ensure the network can be optimized fully, we separate the training process into two phases. In the 1st stage, we only train our network without predicting movable objects in the decoder. Meanwhile, following the previous works [7, 5], we introduce the weighted cross-entropy ($L_{wce}$) and Lovász-Softmax ($L_{ls}$[23] losses to supervise network:
@@
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ mean the weights for losses while $y$ and $\hat{y}$ are the ground truth and predicted results of points. $L(y_{i}^{b},\hat{y}_{i}^{b})$ denote the auxiliary losses for BEV predictions. Moreover, in the 2nd stage, we freeze the pre-trained parameters that are optimized in the 1st stage and only train the rest of network to estimate movable objects by loss function as follows:
@@
Figure 5: Visualization MOS results on SemanticKITTI validation set. Incorrect predictions are marked by blue circles. Best viewed in color and zoom.
@@ IV-A Experimental Settings
Datasets. On SemanticKITTI-MOS [4] dataset and Sipailou-Campus [7] dataset, we compare segmentation performance with previous methods and conduct extensive ablation studies. The SemanticKITTI-MOS dataset is collected by Velodyne HDL-64E LiDAR and contains a total of 22 sequences with labeled point clouds that are remapped from 28 semantic classes into 3 types of motion states. Following the previous algorithms [7, 5, 8], we divide the sequences 00-07, 09-10 (19,130 frames) for training, sequence 08 (4,071 frames) for validation and sequences 11-21 (20,351 frames) for testing. For Sipailou-Campus that is constructed based on solid-state LiDAR, we follow the implementation of [7] to split dataset into 5 train sequences (16,887 frames), 1 validation sequence (3,191 frames) and 2 test sequences (6,201 frames).
@@ IV-C Quantitative Results
TABLE I: Performance comparison on SemanticKITTI validation and test sets. * denotes the method exploiting semantic labels. † means the method is trained both on SemanticKITTI and KITTI-road datasets.
Methods Source IoU (Val) IoU (Test)
@@
Comparison with previous methods. We first evaluate our StreamMOS on SemanticKITTI-MOS benchmark. To ensure fairness, our method is split into two versions in Tab. I, which makes settings as consistent as possible with previous works. Specifically, (a) StreamMOS-V indicates the network that is trained in 1st stage and only uses voxel-based voting as post-processing. (b) StreamMOS-VI* means extra performing 2nd stage training and using instance-based voting that relies on movable object predictions. Then, our methods are compared with existing algorithms, which can be classified as whether semantic annotations are utilized. Specially, the v1 and v2 of MotionSeg3D [5] refer to using kNN or point refine module as post-processing. Moreover, “w/ delay” signifies exploiting point cloud frames within the time window of $[t,t+N]$ to estimate dynamic objects in the $t$ frame. Note that following MotionBEV [7], our results shown in Tab. I are derived from training original SemanticKITTI without any additional data.
TABLE II: Performance comparison on Sipailou-Campus dataset.
Methods Source IoU (Val) IoU (Test)
@@
TABLE III: Comparison of running time (ms) with previous methods.
@@
As illustrated in Tab. I, our streaming method outperforms previous works in most cases. Specifically, our StreamMOS-V exceeds 4DMOS [9] by 1.1% and 7.9% in validation and test. We think that compared with using binary Bayes filter to merge historical results in 4DMOS, our method additionally considers historical feature from the last inference, which can serve as strong spatial priors to improve prediction quality. At the same time, our StreamMOS-VI* surpasses InsMOS* [8] and MF-MOS* [25] in the validation set significantly ($\uparrow$8.4and $\uparrow$5.5%) by instance-based voting. Finally, due to the lack of semantic annotation in Sipailou-Campus dataset, we solely list [4, 7, 5, 9] in Tab. II and confirm the effectiveness of StreamMOS-V, even using a solid-state LiDAR with the narrow field of view and non-repetitive scanning patterns.
Inference Speed. Although our method uses attention mechanism to construct feature association between inferences and merge multiple historical predictions in voting mechanism, it still keeps competitive running time compared with previous approaches in Tab. III. We believe this is the contribution of projection-based backbone, lightweight deformable attention and parameter-free upsampling in decoder, which make our method strike a balance between speed and performance.
@@ IV-D Qualitative Analysis
To compare the previous methods intuitively, we visualize segmentation results in various scenarios. As demonstrated in Fig. 5, LMNet* suffers from boundary-blurring issues in 4th row. Although MotionSeg3D adopts point refinement module to alleviate this problem, it also makes mistakes when dealing with distant objects in the 2nd and 5th rows. Moreover, due to lacking the instance-level sensing, MotionSeg3D is prone to incomplete prediction, such as in the 1st row. Despite adding instance detection like InsMOS* could improve the integrity of segmentation, it will aggravate negative impacts when the prediction is incorrect, as illustrated in the 2nd, 3rd and 5th rows. Unlike the above algorithms, our StreamMOS-VI* first combines observation from multi-view to improve perception of objects at different distances. Then, instead of focusing on a single inference, we build relationship in several inferences to integrate historical feature and predictions, which are used to suppress false results and improve segmentation integrity. Thus, our method can achieve superior performance in Fig. 5.
@@ IV-E Ablation Study
In this section, we conduct comprehensive experiments on the SemanticKITTI validation set to confirm the effectiveness of the proposed modules.
@@
Figure 7: The effect of frame number and BEV size in our SreamMOS-V.
@@
Time Window Length. The time window length determines how long ago predictions can be used by voting mechanism. Thus, we conduct experiments on the time window length to choose the optimal setting for our algorithm. As displayed in Fig. LABEL:fig:time_window_length, the performance will increase rapidly until the length $M$ of time window reaches 8. Despite continuing to raise the length could result in a slight improvement, it requires more time consumption. Thus, we opt for $M$ = 8 as our default.
@@
Furthermore, as shown in Fig. 7(b), compared to previous approaches [8, 9] that require a lot of frames to extract the spatial-temporal features, our method only relies on 3 frames to achieve the best result. We think this is due to the effective reuse of historical feature and predictions in temporal fusion and voting, which has taken rich prior knowledge to network. Meanwhile, too many frames are fed into the network would cause information redundancy and result in degradation.
@@ V Conclusion
In this paper, we analyse the limitations of existing MOS methods and propose a novel streaming structure, which uses memory bank as a bridge to transfer prior information among inferences. Moreover, our StreamMOS captures the complete appearance and motion features of objects from multi-view. To correct false prediction in a single inference, we propose a voting mechanism to integrate historical predictions at voxel and instance levels. The broad experimental results show the effectiveness of proposed modules and prove that our method has competitive performance in diverse aspects.

### v2

@@ III-C Short-term Temporal Fusion
The purpose of this part is to transfer the memory feature $\mathcal{H}_{t-1}$ from the last inference to the present, so that historical spatial states of objects can be retrieved to guide the network in inferring object motion at time $t$. To achieve this, we first build short-term memory bank as a bridge to store $\mathcal{H}_{t-1}$ and connect adjacent inferences. Then, since $\mathcal{F}_{t}$ and $\mathcal{H}_{t-1}$ are not in the same coordinate system, we adopt an attention module with learnable offsets [22] to adaptively find the relationship between two features and combine them by attention weight. Specifically, $\mathcal{H}_{t-1}$ is fed into two linear layers to produce $K$ attention weights $A_{k}$ and sampling offsets $\Delta g_{k}$. Later, based on the offsets $\Delta g_{k}$ and coordinates $g_{k}$ of reference points in $\mathcal{F}_{t}$, a bilinear interpolation is used to gather reference values $G_{k}$ from $\mathcal{F}_{t}$. Finally, $G_{k}$ is weighted by $A_{k}$ to get enhanced feature ^${\hat{\mathcal{H}}_{t}}$. The above process can be formulated as follows:
@@
The $L$ and $K$ are the number of attention heads and reference points, respectively, while $S(\textperiodcentered\cdot\textperiodcentered)$ and $W_{l}$ represent the bilinear sampling and learnable weight of multi-head attention. Next, the ^${\hat{\mathcal{H}}_{t}}$ is processed by normalization layer and feed-forward network (FFN) to generate a updated $\mathcal{H}_{t}$ at the current time:
@@
Here, LN is the layer normalization. Then, $\mathcal{H}_{t}$ is used for two purposes: it replaces $\mathcal{H}_{t-1}$ to update the short-term memory, and it is fed into the decoder to predict segmentation results.
@@ III-D Reduced-parameter Segmentation Decoder
where $\text{head}_{1}$ and $\text{head}_{2}$ consist of several convolution blocks, and $\mathcal{E}_{t}$ is the feature from point-wise encoder. According to discrete classification labels $\mathcal{O}_{t}$, we can acquire the attributes of instance, like location and size, through clustering and use them to optimize $\mathcal{C}_{t}$ in the subsequent voting stage.
Figure 3: Illustration of asymmetric convolution and multi-view features.
@@ III-E Long-term Voting Mechanism
Most existing approaches [4, 5, 19] focus on improving the quality of a single inference through modifying network structure. Nevertheless, in light of the inexplicability and data dependency of neural networks, this strategy may be limited. For example, for a parked car shown in Fig. 1(c), the model may predict it as stationary in one frame and moving in other frames. Meanwhile, due to lacking instance-level perception ability, the network may generate inconsistent results for the different parts of an object, particularly for cars (see Fig. 5).
To solve these problems, we construct a long-term memory bank of length $M$ to store historical predictions and propose a voting module consisting of the voxel-based voting (VBV) and instance-based voting (IBV), which can function as post-processing to correct errors in the current predicted labels $\mathcal{C}_{t}$ using historical results $\{\mathcal{M}_{t-m}\}^{M}_{m=1}$ and movable labels $\mathcal{O}_{t}$. Note that $\mathcal{M}_{t-m}\ (m=1,...,M)$ with coordinates of $\mathcal{P}_{t-m}$ is projected into the coordinate system of $\mathcal{P}_{t}$ to yield $\mathcal{M}^{\prime}_{t-m}$ through the pose transformations $\mathcal{T}_{t-m\rightarrow t}$ in advance.
@@ III-E1 Voxel-based voting
Figure 4: The details of our voting mechanism. It uses voxel-based voting (VBV) and instance-based voting (IBV) to refine coarse predictions.
Motivated by TFNet [23], we first obtain historical predictions $\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1}$ and current forecast $\mathcal{C}_{t}$ in the same coordinate system. Then, we divide points $\mathcal{P}_{t}$ into voxels with fixed size and fill $(\mathcal{C}_{t},\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1})$ into each voxel. Next, as shown in Fig. 4(a), the most frequently predicted label acts as motion state for all points in the same voxel and incorrect labels will be updated. We summarize the above procedure of VBV as: $\Omega(\mathcal{C}_{t},\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1})\mapsto\hat{\mathcal{C}_{t}}$.
@@ III-E2 Instance-based voting
Although VBV can ensure the consistency of motion states within a local area and achieve performance improvement in Tab. IV, it is difficult to achieve instance-level unity, as shown by the output $\hat{\mathcal{C}_{t}}$ from VBV in Fig. 4(b). To solve this, we propose an instance-based voting based on clustering. When given the predicted probability $\mathcal{O}_{t}$ from decoder, we can pick out the foreground points ${\mathcal{P}^{\prime}_{t}}$ from ${\mathcal{P}_{t}}$ and adopt DBSCAN [24] to split ${\mathcal{P}^{\prime}_{t}}$ into $S$ clusters. Then, according to the coordinates of points in each cluster, we can compute $S$ minimum 3D bounding boxes to cover all objects. Thus, we can further crop out instance-level predictions from $\hat{\mathcal{C}_{t}}$ and memory predictions $\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1}$. Finally, similar to voxel-based voting, we adopt the class label with the highest quantity as the motion state for all points in the instance and get the final prediction as: $\Phi(\hat{\mathcal{C}_{t}},\{\mathcal{M}^{\prime}_{t-m}\}^{M}_{m=1},\mathcal{O}_{t})\mapsto{\mathcal{M}_{t}}$.
Finally, when a new refined prediction $\mathcal{M}_{t}$ is output from voting mechanism, we append it to long-term memory while discarding the oldest result $\mathcal{M}_{t-M}$. As a result, compared to relying on the network’s adaptive learning, voting mechanism can explicitly suppress incorrect predictions and improve the consistency of segmentation based on a statistical analysis of historical predictions at the voxel and instance levels.
@@ III-F Loss Functions
To ensure the network can be fully optimized, we separate the training process into two steps. In the first stage, we only train our network without predicting movable objects in the decoder. Meanwhile, following the previous works [7, 5], we introduce the weighted cross-entropy ($L_{wce}$) and Lovász-Softmax ($L_{ls}$[25] losses to supervise network:
@@
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ mean the weights for losses while $y$ and $\hat{y}$ are the ground truth and predicted results of points. $L(y_{i}^{b},\hat{y}_{i}^{b})$ denotes the auxiliary losses for BEV predictions. Moreover, in the second stage, we freeze pre-trained parameters that are optimized in the 1st stage and only train the rest of network to predict movable objects using the following loss function:
@@
Figure 5: The visualization of MOS results on the SemanticKITTI validation set. Incorrect predictions are highlighted, with false negatives marked by green circles and false positives by blue circles. Best viewed in color and zoom.
@@ IV-A Experimental Settings
Datasets. On SemanticKITTI-MOS [4] dataset and Sipailou-Campus [7] dataset, we compare segmentation performance with previous methods and conduct extensive ablation studies. The SemanticKITTI-MOS dataset is collected by a Velodyne HDL-64E LiDAR and contains a total of 22 sequences with labeled point clouds that are remapped from 28 semantic classes into 3 types of motion states. Following the previous algorithms [7, 5, 8], we divide the sequences 00-07, 09-10 for training, sequence 08 for validation and sequences 11-21 for testing. For the Sipailou Campus dataset that is developed on solid-state LiDAR, we follow the implementation of [7] to split dataset into 5 training sequences, 1 validation sequence and 2 test sequences from 26,279 frames.
@@ IV-C Quantitative Results
TABLE I: Performance comparison on SemanticKITTI validation and test sets. * denotes methods that exploit semantic labels. † indicates methods trained on both SemanticKITTI and KITTI-road datasets.
Methods Source IoU (Validation) IoU (Test)
@@
Comparison with Previous Methods. We first evaluate our StreamMOS on SemanticKITTI-MOS benchmark. To ensure fairness, our method is presented in two versions in Tab. I to make settings as consistent as possible with previous works. Specifically, (a) StreamMOS-V indicates the network that is trained in the 1st stage and uses voxel-based voting as post-processing. (b) StreamMOS-VI* means performing extra 2nd stage training and using instance-based voting that relies on movable object predictions. Then, our methods are compared with existing algorithms, which can be classified as whether semantic annotations are utilized. Specially, the two versions of MotionSeg3D [5] refer to using kNN or point refinement as post-processing. Moreover, “w/ delay” signifies exploiting point cloud frames within the time window of $[t,t+N]$ to estimate dynamic objects in the $t$ frame. Note that following MotionBEV [7], our results shown in Tab. I are derived from training original SemanticKITTI without any additional data.
As illustrated in Tab. I, our streaming method outperforms previous works in most cases. Specifically, our StreamMOS-V exceeds 4DMOS [9] by 1.1% and 7.9% in validation and test. We think that compared with using binary Bayes filter to merge historical results in 4DMOS, our method additionally considers historical feature from the last inference, which can serve as strong spatial priors to improve prediction quality. At the same time, our StreamMOS-VI* surpasses InsMOS* [8] and MF-MOS* [27] on the validation set significantly ($\uparrow$8.4and $\uparrow$5.5%) by instance-based voting. Finally, due to the lack of semantic annotation in Sipailou-Campus dataset, we solely list [4, 7, 5, 9] in Tab. II and confirm the effectiveness of StreamMOS-V, even when using a solid-state LiDAR with a narrow field of view and non-repetitive scanning patterns.
TABLE II: Performance comparison on Sipailou Campus dataset.
Methods Source IoU (Validation) IoU (Test)
@@
TABLE III: Comparison of running times (ms) with previous methods.
@@
Inference Speed. Although our method uses attention mechanism to construct feature association between inferences and merge multiple historical predictions in voting mechanism, it still keeps competitive running time compared with previous approaches in Tab. III. We believe this is the contribution of projection-based backbone, lightweight deformable attention and parameter-free upsampling in decoder, which enables our method strike a balance between speed and performance.
@@ IV-D Qualitative Analysis
Advantageous Cases. In Fig. 5, we exhibit the segmentation results in various scenarios to compare the previous methods intuitively. Although MotionSeg3D adopts a point refinement module to alleviate boundary-blurring problem, it still makes mistakes when dealing with distant objects, as shown in the 4th row. Besides, MotionSeg3D tends to produce incomplete segmentation in the 3rd row due to lacking the instance-level sensing. Despite adding instance detection like InsMOS* can improve segmentation integrity, it aggravates negative impact when the prediction is incorrect, as illustrated in the 1st, 2nd and 3rd rows. Unlike these algorithms, our StreamMOS-VI* combines multi-view observations to improve the perception of objects at different distances. Then, we build relationships among several inferences by integrating memory feature and predictions to enhance the segmentation integrity and reduce incorrect results. Thus, we get superior performance in Fig. 5. Failure Cases. Fig. 6 displays that the inaccurate ego poses $\{\mathcal{T}_{t-n\rightarrow t}\}_{n=1}^{N}$ misalign multi-frame point clouds, causing the network to incorrectly infer that the object has moved. While our method mitigates this issue compared to InsMOS*, errors persist as our voting mechanism still relies on precise poses. Thus, we think that developing a MOS model that eliminates the need for pose or implicitly learns ego-motion could be a promising research direction in the future.
@@ IV-E Ablation Study
This part conducts ablation studies on the SemanticKITTI validation set to prove the effectiveness of our method.
@@
Figure 6: A failure case caused by inaccurate ego pose in SemanticKITTI.
@@
Figure 8: The effect of frame number and BEV size on the SreamMOS-V.
@@
Time Window Length. The time window length determines how long ago predictions can be used by voting mechanism. Thus, we conduct experiments on the time window length to choose the optimal setting for our algorithm. As displayed in Fig. 7, the performance will increase rapidly until the length $M$ of time window reaches 8. Despite continuing to raise the length could result in a slight improvement, it requires more time consumption. Thus, we opt for $M$ = 8 as our default.
@@
Furthermore, as shown in Fig. 8(b), compared to previous approaches [8, 9] that require a lot of frames to extract the spatial-temporal features, our method only relies on 3 frames to achieve the best result. We think this is due to the effective reuse of historical feature and predictions in temporal fusion and voting, which bring rich prior knowledge to the network. Meanwhile, feeding too many frames into the network would cause information redundancy and result in degradation.
@@ V Conclusion
In this paper, we analyze the limitations of existing MOS methods and propose a novel streaming structure, which uses memory bank as a bridge to transfer prior information among inferences while capturing the appearance and motion feature of objects from multiple views. To correct false predictions, we propose a voting mechanism to integrate historical results at the voxel and instance levels. Experimental results indicate that our method performs competitively in diverse aspects.